OpenStudy (anonymous):

is this correct?? (6, 3) and (8, 6). Slope (m) = change in y over change in x.? m= y2-y1/x2-x1 my answer >>>> 6 - 8 = -2 3 - 6 = -3 -2/-3 = m (Slope)

hero (hero):

y = mx + b 3 = 6m + b 6 = 8m + b 3 - 6m = b 6 - 8m = b b = b 3 - 6m = 6 - 8m 8m - 6m = 6 - 3 2m = 3 m = 3/2

OpenStudy (anonymous):

how did you get that ??

OpenStudy (anonymous):

thats after you put into point slope correct

hero (hero):

Study it

OpenStudy (anonymous):

?

hero (hero):

It's the same thing, just in a different format.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

is this correct Part 1: 6 - 3 = 3 8 - 6 = 2 3/2 = m (Slope) Part 2: Point Slope Formula = y – y1 = m (x – x1) y - 3 = 3/2 (x - 6)

OpenStudy (anonymous):

why ?

OpenStudy (anonymous):

thats what i have to do for my homework

hero (hero):

Nevermind.

hero (hero):

Continue what you are doing

hero (hero):

By the way, you should remember to post the complete instructions of the problem before posting your answer.

OpenStudy (anonymous):

is it correct thats all i want to know

hero (hero):

How do I know if it's correct if I don't know what the original question is?

hero (hero):

You post the answer, but not the question. No one can help you until you post the question you're working on.

OpenStudy (anonymous):

(6, 3) and (8, 6). Part 1: Slope (m) = change in y over change in x. Part 2: Now find the equation using the point slope formula: Point Slope Formula = y – y1 = m (x – x1) Part 3: Give an example of an equation for a line parallel to your answer from Part 2 Part 4: Give an example of an equation for a line perpendicular to your answer from Part 2